# Leveraging automatic strategy discoveryto teach people how to select better projects

### v1

@@ Abstract
The decisions of individuals and organizations are often suboptimal because normative decision strategies are too demanding in the real world. Recent work suggests that some errors can be prevented by leveraging artificial intelligence to discover and teach prescriptive decision strategies that take people’s constraints into account. So far, this line of research has been limited to simplified decision problems. This article is the first to extend this approach to a real-world decision problem, namely project selection. We develop a computational method (MGPS) that automatically discovers project selection strategies that are optimized for real people and develop an intelligent tutor that teaches the discovered strategies. We evaluated MGPS on a computational benchmark and tested the intelligent tutor in a training experiment with two control conditions. MGPS outperformed a state-of-the-art method and was more computationally efficient. Moreover, the intelligent tutor significantly improved people’s decision strategies. Our results indicate that our method can improve human decision-making in naturalistic settings similar to real-world project selection, a first step towards applying strategy discovery to the real world.
@@ 1 Introduction
Corporations and individuals commonly have to select a project out of multiple alternatives. These project selection problems are usually high-stakes decisions that can be highly impactful for the future of an organization. For example, an organization looking for a sustainable investment project (Khalili-Damghani and Sadi-Nezhad, 2013) can benefit both financially and by improving its public image by selecting an impactful and profitable project, or incur major losses by selecting an unsuitable project.
@@
Previous research on project selection recommends that candidate projects should be evaluated by multiple experts (Coldrick et al., 2002, Khalili-Damghani and Sadi-Nezhad, 2013, Liu et al., 2017), and many structured approaches to integrate the experts’ opinions exist (de Souza et al., 2021). However, existing project selection techniques are not well utilized in the real world (Schmidt and Freeland, 1992, Liu et al., 2017, Henriksen and Traynor, 1999). We hypothesize this underutilization is partly caused by the complexity and implementation costs of normative project selection strategies. Scoring approaches, which score alternative projects on a number of relevant selection criteria, have been proposed as a less complex selection method that can be applied to real-world problems more easily (Henriksen and Traynor, 1999, Khalili-Damghani and Sadi-Nezhad, 2013, Coldrick et al., 2002, Sadi-Nezhad, 2017). However, even scoring approaches face significant challenges when applied to real-world problems: precisely evaluating projects on multiple criteria can be a time-consuming and expensive process, but existing selection techniques often pay little attention to the information-gathering costs and instead assume expert opinions on all criteria are readily available and instead focus on how to integrate existing expert evaluations into a final project choice (Abdel-Basset et al., 2019, Khalili-Damghani and Sadi-Nezhad, 2013).
@@
Due to the limited utilization of complex project selection methods, decision-makers, therefore, often rely on their intuition and much simpler techniques like brainstorming (Kornfeld and Kara, 2013). This is concerning because the intuitive decisions of groups and individuals are highly susceptible to biases and unsystematic errors (Kahneman et al., 1982), such as weighing one’s own opinion too strongly (Yaniv and Kleinberger, 2000), being overconfident, and weighting the recommendations of advisors by their confidence rather than their competence (Bonaccio and Dalal, 2006). However, people’s errors in decision-making can partly be prevented by teaching prescriptive decision strategies adapted to humans by taking their cognitive limitations into account (Lieder and Griffiths, 2020, Callaway et al., 2022). This approach to improving human decision-making is known as boosting (Hertwig and Grüne-Yanoff, 2017).
@@
To identify appropriate decision strategies, we can score candidate strategies by their resource-rationality, that is, the degree to which they make rational use of people’s limited time and bounded cognitive resources (Lieder and Griffiths, 2020). In the resource-rational framework, the decision operations people can perform to arrive at a decision are modeled explicitly and assigned a cost. The overall efficiency of a decision strategy in an environment can then be computed by subtracting the expected costs of the used decision operations from the expected utility of the resulting decision (see Equation 1) (Consul et al., 2022). This measure is called resource-rationality score (RR-score) (Consul et al., 2022). People are usually not fully resource-rational, and identifying decision strategies would enable people to perform as well as possible is an important open problem (Callaway et al., 2022, Consul et al., 2022, Heindrich et al., 2023, Mehta et al., 2022).
@@
Recent work has demonstrated that the theory of resource rationality makes it possible to leverage AI to automatically discover and teach decision strategies that enable real people to make their decisions as well as possible (Callaway et al., 2022, Consul et al., 2022, Becker et al., 2022, Skirzyński et al., 2021, Mehta et al., 2022). This approach has been dubbed AI-powered boosting. The first step of AI-powered boosting is to compute resource-rational decision strategies. Automatic strategy discovery methods (Callaway et al., 2022, Consul et al., 2022, Heindrich et al., 2023, Skirzyński et al., 2021, Mehta et al., 2022) can discover efficient decision strategies by solving the metareasoning problem of deciding which decision operations to perform. While recent work has extended automatic strategy discovery methods to larger (Consul et al., 2022) and partially observable environments (Heindrich et al., 2023), so far, they have not been applied to real-world decisions such as project selection.
@@
Developing computational methods for decision support is one of the core problems of operations research (Gupta et al., 2022, Eom and Kim, 2006). In this article, we present a novel decision support method with the goal of discovering prescriptive decision-strategies that can be taught to people to improve how they select projects. Project selection is challenging because many crucial attributes of the candidate projects, such as their expected profitability, cannot be observed directly. Instead, they have to be inferred from observations that are not fully reliable. We, therefore, formalize project selection strategies as policies for solving a particular class of partially observable Markov Decision Processes (POMDPs). This formulation allows us to develop the first algorithm for discovering resource-rational strategies for human project selection. To achieve this, we model a realistic project selection task as a metareasoning problem. The metareasoning consists in deciding which information one should request from which advisors when information is highly limited, uncertain, and costly. We develop an efficient algorithm for solving this problem based on MGPO, an algorithm for discovering resource-rational strategies for human decision-making in partially observable environments (Heindrich et al., 2023), and apply it to derive a prescriptive decision strategy for a project selection problem a financial institution faced in the real world (Khalili-Damghani and Sadi-Nezhad, 2013). Finally, we develop an intelligent tutor (Callaway et al., 2022) that teaches the decision strategy discovered by our method to people. We evaluated our approach by letting our intelligent tutor teach the automatically discovered project selection strategy to about 100 people, and then evaluated the quality of their decisions in realistic project selection problems against two control groups. Our results indicate that our approach can successfully improve human decisions in real-world problems where people are reluctant to let machines decide for them.
@@ 2 Background
Resource-rational analysis is a cognitive modeling paradigm used to describe optimal decision-making with limited computational resources (Lieder and Griffiths, 2020). The paradigm considers not only the quality of decisions, but also the computational costs of reaching them. According to this paradigm, good decision-making consists in making efficient use of the available computational resources by balancing the competing objectives of maximizing the expected decision quality and minimizing computational costs. Resource rationality can be used to describe the quality of a decision strategy by subtracting the cost of the cognitive operations from the expected reward of the resulting decision. This measure is called resource-rationality score (RR-score) (Consul et al., 2022). It will be used throughout this article to evaluate and compare alternative decision strategies. People are usually not fully resource-rational, and figuring out what decision strategies would enable people to perform as well as possible is an important open problem (Callaway et al., 2022, Consul et al., 2022, Heindrich et al., 2023, Mehta et al., 2022).
@@
Discovering resource-rational planning strategies can be achieved by solving a meta-level Markov decision process (Hay et al., 2014, Callaway et al., 2018a, b), which models the metareasoning process as a Markov Decision Process (MDP), which state represents the current belief about the environment and actions represent decision operations. Performing a decision operation results in a negative cost and results in an update to the belief state. A special termination action represents exiting the metareasoning process and making a real-world decision, guided by the current beliefs (Hay et al., 2014). Multiple methods for solving meta-level MDPs exist (e.g. (Russell and Wefald, 1991, Callaway et al., 2018a, Hay et al., 2014, Consul et al., 2022)). We refer to these algorithms as strategy discovery methods (Callaway et al., 2018a, 2022, Skirzyński et al., 2021, Consul et al., 2022, Heindrich et al., 2023, Mehta et al., 2022). They learn or compute policies for selecting sequences of cognitive operations (i.e., computations) people can perform to reach good decisions.
MGPO (Heindrich et al., 2023) is currently the only strategy discovery algorithm that can efficiently approximate resource-rational strategies for decision-making in partially observable environments. MGPO chooses decision operations by approximating their value of computation (Russell and Wefald, 1991) in a myopic manner: it always selects the computation whose execution would yield the highest expected gain in reward if a decision had to be made immediately afterward, without any additional planning. MGPO makes use of the myopic approximation due to the high computational complexity of partially observable Markov Decision Processes (POMDP). Alternative approaches to modeling information gathering POMDPs are difficult to apply to partially observable meta-level MDPs, as they require carefully designed information rewards (Boutilier, 2002) or simplify the problem by restricting environments to be symmetric (Doshi and Roy, 2008).
@@
Past work has developed cognitive tutors that teach automatically discovered planning strategies to people (Callaway et al., 2022, Skirzyński et al., 2021, Mehta et al., 2022, Consul et al., 2022). Training experiments indicated that training with these cognitive tutors could significantly boost the quality of people’s planning and decision-making (Callaway et al., 2022, Skirzyński et al., 2021, Consul et al., 2022, Heindrich et al., 2023, Mehta et al., 2022). These cognitive tutors teach efficient decision strategies in an automated manner, usually by computing the value of available decision operations using strategy discovery methods, and providing the learner feedback on the quality of the computations they select. Initially limited to small planning tasks due to the computational complexity of solving meta-level MDPs (Lieder et al., 2019, Callaway et al., 2022), recent work has extended existing methods to larger (Consul et al., 2022) and partially observable (Heindrich et al., 2023) settings. However, none of these methods have been applied to naturalistic problems so far.
@@
In this section, we introduce explain our general resource-rational model of project selection, which we expect to be widely applicable to concrete, real-world project selection problems.
@@
Our model of project selection consists of two decision problems, an object-level decision-problem and a meta-level MDP (Callaway et al., 2018a, Hay et al., 2014). The two decision problems separate the actions the decision-maker has to choose between (object level), such as executing one project versus another, from decision operations that represent thinking about which of those object-level actions to perform (meta-level), such as gathering information about the projects’ attributes. This allows us to solve both problems separately. The object-level decision problem is a MCDM problem, where a set of potential projects are evaluated using relevant criteria weighted by fixed predetermined weights . Actions in the object-level problem represent selecting the corresponding project (). The reward of a selecting a project is computed by summing the weighted criteria scores of the selected project: (Coldrick et al., 2002).
@@
While the object-level decision problem is our model of the project selection task, the meta-level MDP is our formalization of the problem of discovering resource-rational project selection strategies. It models the task of gathering information about deciding which project to select. States in the meta-level MDP are belief states that represent the current information about each project’s attributes. We model belief states using a multivariate Normal distribution to quantify the estimated value and uncertainty about the projects’ scores on the criteria: . The actions (decision operations) of the meta-level MDP gather information about the different attributes of projects by asking one of the experts for their estimate of how one of the projects scores on one of the criteria. Experts provide discrete estimates from to , and expert estimates can differ in their reliability and their cost. Specifically, the available actions are , where the meta-level action represents asking the expert for their estimate of criterion of project . After receiving information from an expert, the current belief state is updated by applying the update equation for a Gaussian likelihood function with standard deviation (i.e. the expert’s reliability) and a conjugate Gaussian prior (i.e., the current belief), that is and .
@@
The reward of these meta-level actions is the negative cost of asking the expert for help. Finally, the meta-level action is the termination action, which corresponds to terminating the decision-making process and selecting a project. The reward of the termination action is the expected reward of selecting the best project according to the current belief-state. An optional budget parameter specifies the maximum number of available meta-level actions, after which the termination action is performed automatically.
@@
Meta-level MDPs are notoriously difficult to solve due to their extremely large state space (Consul et al., 2022, Hay et al., 2014). In the project selection task, the meta-level MDP has possible belief states and possible meta-level actions. Our meta-level MDP introduces multiple new intricacies that current metareasoning methods like MGPO (Heindrich et al., 2023) aren’t equipped to handle, making strategy discovery in this setting especially difficult. Compared to previously formulated meta-level MDPs (Callaway et al., 2018a, Hay et al., 2014, Consul et al., 2022, Heindrich et al., 2023, Mehta et al., 2022), our meta-level description of project selection differs in the following ways: (1) the maximum amount of meta-level actions is constrained with a budget, (2) the project selection task features multiple consultants who differ in both the quality of their advice and the cost of their services, (3) consultants in the project selection task offer discrete estimates of each criterion , requiring that (4) criteria ratings are scaled to allow weighting the criteria according to their importance.

### v2

# Discovering the curriculum with AI: A proof-of-concept demonstration with an intelligent tutoring system

@@ Abstract
The decisions of individuals and organizations are often suboptimal because fully rational decision-making is too demanding in the real world. Recent work suggests that some errors can be prevented by leveraging artificial intelligence to discover and teach clever heuristics. So far, this line of research has been limited to simplified, artificial decision-making tasks. This article is the first to extend this approach to a real-world decision problem, namely, executives deciding which project their organization should launch next. We develop a computational method (MGPS) that automatically discovers project selection strategies that are optimized for real people, and we develop an intelligent tutor that teaches the discovered project selection procedures. We evaluated MGPS on a computational benchmark and tested the intelligent tutor in a training experiment with two control conditions. MGPS outperformed a state-of-the-art method and was more computationally efficient. Moreover, people who practiced with our intelligent tutor learned significantly better project selection strategies than the control groups. These findings suggest that AI could be used to automate the process of discovering and formalizing the cognitive strategies taught by intelligent tutoring systems.
@@ 1 Introduction
Teaching general cognitive skills is an essential goal of education (Collins et al., 1991). One way to represent cognitive skills is through cognitive strategies. For example, when teaching how to effectively solve math problems, instructors might teach their students metacognitive strategties such as evaluating the effectiveness of different solution methods or breaking problems into smaller subproblems. Prior work has shown that teaching students skills such as carefully evaluating their problem understanding of math problems can significantly improve their test scores (Özsoy and Ataman, 2009). Similarly, students’ writing quality can be improved by teaching them strategies for planning and revising their texts (Graham and Perin, 2007).
The teaching of cognitive strategies can be automated by developing intelligent tutoring systems (Anderson et al., 1985, Koedinger et al., 1997, Graesser et al., 2012). In current applications, the strategies the intelligent tutoring systems teach have to be specified manually (Aleven et al., 2009, Chi and VanLehn, 2010, Özsoy and Ataman, 2009). While hand-crafted strategies can often be effective, they rely on the designer’s intuitions. This constitutes a key bottleneck, as it presumes that designers know the optimal cognitive strategies and can precisely articulate them. While this may be the case in some limited domains, in many domains it is unknown which cognitive strategies are optimal. Expertise in areas such as mathematical problem solving often involves tacit, procedural knowledge that is difficult to verbalize (Anderson, 1982). One way to overcome this bottleneck is to develop rational process models of human cognition (Lieder and Griffiths, 2020). Recent technical advances have made it possible to automate this process by leveraging AI to discover near-optimal cognitive strategies automatically (Callaway et al., 2022a, Krueger et al., 2024). Adopting the strategies discovered by these methods consistently allowed people to perform better than the strategies they intuitively used (Callaway et al., 2022a, Consul et al., 2022, Heindrich et al., 2025). Moreover, it is now possible to develop intelligent tutotring systems that use AI to discover the cognitive strategies they teach by themselves (Callaway et al., 2022a). However, so far, this approach has only been evaluated in artificial decision-making tasks. Therefore, its viability for educational applications is still unclear.
In this article, we present the first proof-of-concept that automatic strategy discovery can inform educational applications in the real world. We build directly on our prior work (Heindrich et al., 2025, Consul et al., 2022), which brought strategy discovery closer to real-world application by introducing a metareasoning model of planning in partially observable environments. Although the planning task was complex, it was still highly artificial and unrealistic. Here, we show that the general approach to combine automatic strategy discovery with an intelligent tutor can be extended to an important class of real-world decisions: deciding which project an organization, team, or individual should pursue next. This class of decision problems is known as project selection.
@@
Corporations and individuals often need to choose a project from multiple alternatives. These project selection problems are usually high-stakes decisions that can be highly impactful for the future of an organization. For example, an organization looking for a sustainable investment project (Khalili-Damghani and Sadi-Nezhad, 2013) can benefit both financially and by improving its public image by selecting an impactful and profitable project, or incur major losses by selecting an unsuitable project.
@@
Previous research on project selection recommends that candidate projects should be evaluated by multiple experts (Coldrick et al., 2002, Khalili-Damghani and Sadi-Nezhad, 2013, Liu et al., 2017), and many structured approaches to integrate the experts’ opinions exist (de Souza et al., 2021). However, existing project selection techniques are not well utilized in the real world (Schmidt and Freeland, 1992, Liu et al., 2017, Henriksen and Traynor, 1999). We hypothesize this underutilization is partly caused by the complexity and implementation costs of normative project selection strategies. Scoring approaches, which score alternative projects on several relevant selection criteria, have been proposed as a less complex selection method that can be applied to real-world problems more easily (Henriksen and Traynor, 1999, Khalili-Damghani and Sadi-Nezhad, 2013, Coldrick et al., 2002, Sadi-Nezhad, 2017). However, even scoring approaches face significant challenges when applied to real-world problems: precisely evaluating projects on multiple criteria can be a time-consuming and expensive process, but existing selection techniques often pay little attention to the information-gathering costs and instead assume expert opinions on all criteria are readily available and instead focus on how to integrate existing expert evaluations into a final project choice (Abdel-Basset et al., 2019, Khalili-Damghani and Sadi-Nezhad, 2013).
@@
Due to the limited utilization of complex project selection methods, decision-makers, therefore, often rely on their intuition and much simpler techniques like brainstorming (Kornfeld and Kara, 2013). This is concerning because the intuitive decisions of groups and individuals are highly susceptible to biases and unsystematic errors (Kahneman et al., 1982), such as weighing one’s own opinion too strongly (Yaniv and Kleinberger, 2000), being overconfident, and weighting the recommendations of advisors by their confidence rather than their competence (Bonaccio and Dalal, 2006). However, people’s errors in decision-making can partly be prevented by teaching prescriptive decision strategies adapted to humans by taking their cognitive limitations into account (Lieder and Griffiths, 2020, Callaway et al., 2022a). This approach to improving human decision-making is known as boosting (Hertwig and Grüne-Yanoff, 2017).
@@
To identify appropriate decision strategies, we can score candidate strategies by their resource-rationality, that is, the degree to which they make rational use of people’s limited time and bounded cognitive resources (Lieder and Griffiths, 2020). In the resource-rational framework, the decision operations people can perform to arrive at a decision are modeled explicitly and assigned a cost. The overall efficiency of a decision strategy in an environment can then be computed by subtracting the expected costs of the used decision operations from the expected utility of the resulting decision (see Equation 1) (Consul et al., 2022). This measure is called resource-rationality score (RR-score) (Consul et al., 2022). People are usually not fully resource-rational, and identifying decision strategies would enable people to perform as well as possible is an important open problem (Callaway et al., 2022a, Consul et al., 2022, Heindrich et al., 2025, Mehta et al., 2022).
@@
Recent work has demonstrated that the theory of resource rationality makes it possible to leverage AI to automatically discover and teach decision strategies that enable real people to make their decisions as well as possible (Callaway et al., 2022a, Consul et al., 2022, Becker et al., 2022, Skirzyński et al., 2021, Mehta et al., 2022). This approach has been dubbed AI-powered boosting. The first step of AI-powered boosting is to compute resource-rational decision strategies. Automatic strategy discovery methods (Callaway et al., 2022a, Consul et al., 2022, Heindrich et al., 2025, Skirzyński et al., 2021, Mehta et al., 2022) can discover efficient decision strategies by solving the metareasoning problem of deciding which decision operations to perform. While recent work has extended automatic strategy discovery methods to larger (Consul et al., 2022) and partially observable environments (Heindrich et al., 2025), so far, they have not been applied to real-world decisions such as project selection.
@@
In this article, we present a novel decision support method with the goal of discovering prescriptive decision-strategies that can be taught to people to improve how they select projects. Project selection is challenging because many crucial attributes of the candidate projects, such as their expected profitability, cannot be observed directly. Instead, they have to be inferred from observations that are not fully reliable. We, therefore, formalize project selection strategies as policies for solving a particular class of partially observable Markov Decision Processes (POMDPs). This formulation allows us to develop the first algorithm for discovering resource-rational strategies for human project selection. To achieve this, we model a realistic project selection task as a metareasoning problem. The metareasoning consists in deciding which information one should request from which advisors when information is highly limited, uncertain, and costly. We develop an efficient algorithm for solving this problem based on MGPO, an algorithm for discovering resource-rational strategies for human decision-making in partially observable environments (Heindrich et al., 2025), and apply it to derive a prescriptive decision strategy for a project selection problem a financial institution faced in the real world (Khalili-Damghani and Sadi-Nezhad, 2013). Finally, we develop an intelligent tutor (Callaway et al., 2022a) that teaches the decision strategy discovered by our method to people. We evaluated our approach by letting our intelligent tutor teach the automatically discovered project selection strategy to about 100 people, and then evaluated the quality of their decisions in realistic project selection problems against two control groups. Our results indicate that our approach can successfully improve human decisions in real-world problems where people are reluctant to let machines decide for them.
@@ 2 Background
Resource-rational analysis is a cognitive modeling paradigm used to describe optimal decision-making with limited computational resources (Lieder and Griffiths, 2020). The paradigm considers not only the quality of decisions, but also the computational costs of reaching them. According to this paradigm, good decision-making consists in making efficient use of the available computational resources by balancing the competing objectives of maximizing the expected decision quality and minimizing computational costs. Resource rationality can be used to describe the quality of a decision strategy by subtracting the cost of the cognitive operations from the expected reward of the resulting decision. This measure is called resource-rationality score (RR-score) (Consul et al., 2022). It will be used throughout this article to evaluate and compare alternative decision strategies. People are usually not fully resource-rational, and figuring out what decision strategies would enable people to perform as well as possible is an important open problem (Callaway et al., 2022a, Consul et al., 2022, Heindrich et al., 2025, Mehta et al., 2022).
@@
Discovering resource-rational planning strategies can be achieved by solving a meta-level Markov decision process (Hay et al., 2014, Callaway et al., 2018a, b), which models the metareasoning process as a Markov Decision Process (MDP), which state represents the current belief about the environment and actions represent decision operations. Performing a decision operation results in a negative cost and results in an update to the belief state. A special termination action represents exiting the metareasoning process and making a real-world decision, guided by the current beliefs (Hay et al., 2014). Multiple methods for solving meta-level MDPs exist (e.g. (Russell and Wefald, 1991, Callaway et al., 2018a, Hay et al., 2014, Consul et al., 2022)). We refer to these algorithms as strategy discovery methods (Callaway et al., 2018a, 2022a, Skirzyński et al., 2021, Consul et al., 2022, Heindrich et al., 2025, Mehta et al., 2022). They learn or compute policies for selecting sequences of cognitive operations (i.e., computations) people can perform to reach good decisions.
MGPO (Heindrich et al., 2025) is currently the only strategy discovery algorithm that can efficiently approximate resource-rational strategies for decision-making in partially observable environments. MGPO chooses decision operations by approximating their value of computation (Russell and Wefald, 1991) in a myopic manner: it always selects the computation whose execution would yield the highest expected gain in reward if a decision had to be made immediately afterward, without any additional planning. MGPO makes use of the myopic approximation due to the high computational complexity of partially observable Markov Decision Processes (POMDP). Alternative approaches to modeling information gathering POMDPs are difficult to apply to partially observable meta-level MDPs, as they require carefully designed information rewards (Boutilier, 2002) or simplify the problem by restricting environments to be symmetric (Doshi and Roy, 2008).
Intelligent tutoring systems
Intelligent tutoring systems (ITS) are automated computer systems that teach skills in digital learning environments (Graesser et al., 2012). They provide a highly scalable learning environment (Koedinger et al., 1997) that often can be similarly effective or even more effective than human tutors (VanLehn, 2011). ITS achieve this by offering a personalized learning environment in which learners are able to practice skills interactively, often supported through a cognitive model of the learner’s progress that allows to adapt the learning process to their current needs (Woolf, 2010). Recently, ITS have also been applied to teach general cognitive skills such as help-seeking (Aleven et al., 2006) or reading comprehension (Guerra and Mellado, 2017).
@@
Past work has developed cognitive tutors that teach automatically discovered planning strategies to people (Callaway et al., 2022a, Skirzyński et al., 2021, Mehta et al., 2022, Consul et al., 2022). Training experiments indicated that training with these cognitive tutors could significantly boost the quality of people’s planning and decision-making (Callaway et al., 2022a, Skirzyński et al., 2021, Consul et al., 2022, Heindrich et al., 2025, Mehta et al., 2022). These cognitive tutors teach efficient decision strategies in an automated manner, usually by computing the value of available decision operations using strategy discovery methods, and providing the learner feedback on the quality of the computations they select. Initially limited to small planning tasks due to the computational complexity of solving meta-level MDPs (Lieder et al., 2019, Callaway et al., 2022a), recent work has extended existing methods to larger (Consul et al., 2022) and partially observable (Heindrich et al., 2025) settings. However, none of these methods have been applied to naturalistic problems so far.
@@
In this section, we introduce our general resource-rational model of project selection, which we expect to be widely applicable to concrete, real-world project selection problems.
@@
Our model of project selection consists of two decision problems, an object-level decision-problem and a meta-level MDP (Callaway et al., 2018a, Hay et al., 2014). The two decision problems separate the actions the decision-maker has to choose between (object level), such as executing one project versus another, from decision operations that represent thinking about which of those object-level actions to perform (meta-level), such as gathering information about the projects’ attributes. This allows us to solve both problems separately. The object-level decision problem is a MCDM problem, where a set of potential projects are evaluated using relevant criteria weighted by fixed predetermined weights . Actions in the object-level problem represent selecting the corresponding project (). The reward of selecting a project is computed by summing the weighted criteria scores of the selected project: (Coldrick et al., 2002).
While the object-level decision is to select a project, the meta-level MDP is our formalization of the problem of discovering resource-rational project selection strategies. A project selection strategy is a systematic procedure for gathering decision-relevant information about various projects and turning it into a decision. We model this information gathering as a sequence of decisions about whether to request additional information and, if so, which information to request next. In each step, the decision-maker can ask one expert to rate one feature of one project. They can then use the resulting information to decide which, if any, additional information to request in the next step. Different project selection strategies differ in 1) which information they request depending on what is already known, and 2) when they stop gathering information and select a project based on what is already known. Resource-rational project selection strategies make these choices near optimally to achieve the best possible trade-off between the quality of the selected project and the cost of the time and information invested to select it.
@@
States in the meta-level MDP are belief states that represent the current information about each project’s attributes. We model belief states using a multivariate Normal distribution to quantify the estimated value and uncertainty about the projects’ scores on the criteria: . The actions (decision operations) of the meta-level MDP gather information about the different attributes of projects by asking one of the experts for their estimate of how one of the projects scores on one of the criteria. Experts provide discrete estimates from to , and expert estimates can differ in their reliability and their cost. Specifically, the available actions are , where the meta-level action represents asking the expert for their estimate of criterion of project .
Equations 2 and 3 describe the belief update. After receiving information from an expert, the current belief state is updated by applying the conjugate Gaussian update, which integrates the new observation with standard deviation (the expert’s reliability) into the current belief. Beliefs are scaled by their respective criteria weight throughout.
@@
The reward of these meta-level actions is the negative cost of asking the expert for help. Finally, the meta-level action is the termination action, which corresponds to terminating the decision-making process and selecting a project. The reward of the termination action is the expected reward of selecting the best project according to the current belief state. An optional budget parameter specifies the maximum number of meta-level actions that can be performed before the termination operation is selected.
@@
Meta-level MDPs are notoriously difficult to solve due to their extremely large state space (Consul et al., 2022, Hay et al., 2014). In the project selection task, the meta-level MDP has possible belief states and possible meta-level actions. Our meta-level MDP introduces multiple new intricacies that current metareasoning methods like MGPO (Heindrich et al., 2025) aren’t equipped to handle, making strategy discovery in this setting especially difficult. Compared to previously formulated meta-level MDPs (Callaway et al., 2018a, Hay et al., 2014, Consul et al., 2022, Heindrich et al., 2025, Mehta et al., 2022), our meta-level description of project selection differs in the following ways: (1) the maximum amount of meta-level actions is constrained with a budget, (2) the project selection task features multiple consultants who differ in both the quality of their advice and the cost of their services, (3) consultants in the project selection task offer discrete estimates of each criterion, requiring that (4) criteria ratings are scaled to allow weighting the criteria according to their importance.